kaylie:

After Gwen, Tristan, and Keith finish exercising, they go to the fair. At the fair, they each pay the entry fee and also buy tickets they can use for food or rides. Gwen pays the entry fee and buys 10 tickets. It costs her a total of $30. Tristan pays the entry fee and buys 15 tickets. It costs him a total of $40. Keith pays the entry fee and buys 10 tickets. It costs him a total of $30. In this task, you will create a system of equations and find the cost of each ticket. Let x represent the entry fee and y represent the cost of each ticket in dollars. Part A Write an equation representing how much Gwen paid.

supie:

The equation will be written as \[x + 10y = 30\]\[x + 15y = 40 \]

supie:

Then we subtract those two The bigger one goes on the top \[x+15y=40\] \[x+10y=30\] So, \[15y-10y=10\] and \[40-30=10\]

supie:

\(\color{#0cb}{\text{Originally Posted by}}\) @supie Then we subtract those two The bigger one goes on the top \[x+15y=40\] \[x+10y=30\] So, \[15y-10y=10\] and \[40-30=10\] \(\color{#0cb}{\text{End of Quote}}\) So now we have \[5y = 10\] And \[10/5\] or \[10รท5\]\[=?\]